after a lot of test it turned out that the optically contacted QWP/PBS combinations used for the reference cavity so far are very bad alligned.
We tested two out of three we have in the PSL lab and both are bad, meaning about 10% of the linear polarized light entering the PBS are not converted into circular polarized light and so not reflected when comming in the reverse direction. By replacing the optically contacted version by individual PBS and QWP the amount of wrong light dropped by a factor of 100 or so.
Replacing the bad optics should reduce the effect from backscattered/reflected light, which increases the RIN a lot at low frequencies. It doesn't seem to be the laser itself, as with none, one or two FI the spectrum seems to be the same bad level when light is reflected back into the PMC. It looks like the source is the PMC itself or it's control loop.
So, Tara is replacing the bad ones and re-aligning everything. Temp is good, both cavities are resonant at the same time so i hope we can get new/better data tomorrow.